A shop assistant makes 3 entry errors per week. Find the Probability that in a week chosen at random she will make:



(i) Exactly 2 errors [2 marks]



(ii) At least 2 errors [3 marks]



(b) What is the probability that over a period of two weeks, the shop assistant in part (b) above will commit at least 3 errors?

1
Expert's answer
2022-06-02T15:31:20-0400

a) "\lambda=3"

(i)

"P(X=2)=\dfrac{e^{-3}(3)^2}{2!}=0.224042"

(ii)

"P(X\ge2)=1-P(X=0)-P(X=1)""=1-\dfrac{e^{-3}(3)^0}{0!}-\dfrac{e^{-3}(3)^1}{1!}=0.800852"

b) "\lambda t=3(2)=6"


"P(X\ge3)=1-P(X=0)"




"-P(X=1)-P(X=2)"

"=1-\dfrac{e^{-6}(6)^0}{0!}-\dfrac{e^{-6}(6)^1}{1!}-\dfrac{e^{-6}(6)^2}{2!}"

"=0.938031"